OpenStudy (anonymous):

please help me on limits. 1.) lim x -> 2 sec(piex/6) 2.) Lim -> 0 [square root(2+x)- square root (2)]/ x 3.) lim -> 1 (lnx)/x

OpenStudy (johnweldon1993):

1) \(\large \lim_{x \rightarrow 2} sec(\frac{\pi x}{6}) = sec(\frac{2\pi}{6}) = sec(\frac{\pi}{3}) = \frac{1}{cos(\frac{\pi}{3})} =?\)

OpenStudy (johnweldon1993):

\[\large \lim_{x \rightarrow 0} \frac{\sqrt{2 + x} - \sqrt{2}}{x}\] For this...have you learned L'Hospital yet?

jimthompson5910 (jim_thompson5910):

for #2, you can multiply top and bottom by \(\LARGE \sqrt{2+x}+\sqrt{2}\) to rationalize the numerator

OpenStudy (anonymous):

@johnweldon1993 no, I haven't.